OpenStudy (mylahb):

x^2+6x-3=0 Solve the equation by completing the square. @cj49

OpenStudy (cj49):

here it wud be x^2+6x=-3

OpenStudy (mylahb):

so it would be x^2+6x+3=-3+3?

OpenStudy (cj49):

half of 6 is 3 then ull needoth the sides to square it 3x3=9 and then add it to b